OpenStudy (anonymous):

Using the ratio of perfect squares method, what is square root of 128 rounded to the nearest hundredth?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@caitesinclair @KyanTheDoodle @inowalst @OpenStudyRocks5*

OpenStudy (michele_laino):

we can write: \[128 = 2 \times {8^2}\]

OpenStudy (anonymous):

right

OpenStudy (michele_laino):

so, we have: \[\sqrt {128} = \sqrt {2 \times {8^2}} = 8\sqrt 2 = ...?\]

OpenStudy (anonymous):

8 x the square root of 2?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

\[\large \sqrt {128} = \sqrt {2 \times {8^2}} = 8 \times \sqrt 2 = ...?\]

OpenStudy (anonymous):

1.2 is the square root of 2 right

OpenStudy (michele_laino):

better is: \[\sqrt 2 = 1.414\]

OpenStudy (anonymous):

11.31?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thanks but where did you get 8 from ? when we started